OpenStudy (princessaurora):

*Trigonometric Functions* Consider the equation 3tan2(x)−1=0. The solutions of this equation over the interval [0, π] are? *Please help will give medal*

OpenStudy (anonymous):

solve \[\tan(x)=\pm\frac{1}{\sqrt{3}}\]

OpenStudy (princessaurora):

x=π6±πn,7π6±πn

OpenStudy (princessaurora):

Is that correct?

OpenStudy (anonymous):

yeah i guess \(\frac{\pi}{6}\) works for sure

OpenStudy (anonymous):

yeah that looks good except it said only on the interval \[[0,\pi]\]

OpenStudy (princessaurora):

what about the second one?

OpenStudy (anonymous):

you have a general solution there you only want the solution on \([0,\pi]\) not a general one

OpenStudy (princessaurora):

ok so can you help me with getting that?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\frac{\pi}{6}\] works because \[\tan(\frac{\pi}{6})=\frac{1}{\sqrt3}\] do you know why (just asking)

OpenStudy (princessaurora):

because you subtracted it?

OpenStudy (anonymous):

the other one on that interval is \(\frac{5\pi}{6}\)

OpenStudy (princessaurora):

Okay thanks so much! I appreciate it I got it now.

OpenStudy (anonymous):

yw